OpenStudy (anonymous):

The age of father 8 years back was 5 times that of his son. After 8 years, his age will be 8 years more than the double the age of his son. Find their present ages. (Use Elimination by Substitution method to solve)

OpenStudy (anonymous):

F - 8 = 5(S - 8) and F + 8 = 2(S + 8) + 8 F - 5S = -32 and F - 2S = 16 Solving by elimination: -3S = -48 -> Son is 16, so Father is 48

OpenStudy (anonymous):

Thanks!! Cheers. :D

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @bajji